OpenStudy (anonymous):

Does the point (2, -4) lie on the line with equation x + 2y + 4 = 0? Explain.

OpenStudy (anonymous):

1(2) + 2(-4) + 4 = 0 2 - 8 + 4 = 0 -6 + 4 = 0 -2 = 0? NO

OpenStudy (anonymous):

To clarify what pk51 wrote, you start out by inputting the x and y values given into the original equation: This gives you 2 + 2(-4)+4=0 This becomes 2 - 8 + 4 = 0 Which is -2 = 0. This is an impossibility, as -2 is not equal to zero, and if the point were on the line, then -2 = -2 or 0 = 0 Hope this helps. If it does, please give me a medal. Visit my website at: http://www.tutorsean.net for free math videos.

OpenStudy (anonymous):

Thankyou so much sean. You were a great help.